Tuesday 7th December 2021

IP-Only Fibernät Automatiskt uppdatering: NORRTÄLJE

Ticket Number: INC2003733
Business Area: DATACOMMUNICATIONS
Status: RESOLVED
Fault Type: OUTAGE
Location: NORRTĂ„LJE
Occured: 07/12/2021 15:50 UTC
Estimated Time of Repair: 09/12/2021 00:00 UTC
Service(s) Restored: 08/12/2021 21:30 UTC

Details:
07/12/2021 19:17 UTC Increased affected by disturbance. FE is on site of suspected cause for disturbance for further troubleshooting. Several other FE are going to other sites to perform measurements. No ETR available.
07/12/2021 20:36 UTC Technician measured location of potential cable cut. Technician is on his way to locate fault.
07/12/2021 22:01 UTC Cable fault detected. Technicians are working on remedying the issue.
08/12/2021 01:00 UTC Repair work is ongoing.
08/12/2021 02:07 UTC Repair work is ongoing.
08/12/2021 03:20 UTC Repair work is ongoing.
08/12/2021 04:47 UTC The estimated time of repair is pushed back because the fault was more substantial than initial estimates.
Due to prevailing circumstances of nature and weather.
08/12/2021 07:23 UTC Repair work is ongoing. ETR changed due to prevailing circumstances of nature and weather.
08/12/2021 08:04 UTC Repair work is ongoing.
Splicebox damaged by ice due to water leakage. Work is underway to heat up the splicebox so that technicians can repair the broken cable.
08/12/2021 08:38 UTC Repair work is ongoing.
08/12/2021 09:32 UTC Repair work is ongoing.
08/12/2021 09:54 UTC Repair work is ongoing.
The ice has been removed from the splicebox and technicians are preparing to start repairing the cable.
08/12/2021 10:53 UTC Repair work is ongoing.
08/12/2021 12:15 UTC Repair work is ongoing.
08/12/2021 15:03 UTC Repairs are still on-going. Due to the cold weather and the damage to the splicing box repairs are progressing much slower than estimated. We are investigating the possibility of a second team preparing work in parallel to have as much ready as possible when splicing can begin.
08/12/2021 16:56 UTC In case it will be needed, a second team is collecting materials for blowing new cable in parallell with the ongoing repairs to the splicebox.
08/12/2021 19:27 UTC Repair work is ongoing.
08/12/2021 21:18 UTC Some circuits have been restored. Splicing continues.
08/12/2021 22:04 UTC Repair work is ongoing.
Some circuits have been restored. Splicing continues.
08/12/2021 23:33 UTC Outage is resolved and services are restored after the cable fault was fixed. If you still experience any issues please contact GlobalConnect Service Desk.